### v1

@@ 3. $(iii)$
    If a sub-$S$-algebra $S^{\prime}$ of $S^{\sigma}$ is $(S_{2})$, then $S^{\prime}=S^{\sigma}$.
@@
Let $S^{\prime}$ be a sub-$S$-algebra of $R^{\nu}$. Let $\mathfrak{q}^{\prime}\in\operatorname{Spec}_{1}(S^{\prime})$. Then $\operatorname{ht}(\mathfrak{q}^{\prime}\cap R)=1$ by Lemma 4.10, so $\operatorname{ht}(\mathfrak{q}^{\prime}\cap S)=1$. Therefore $S^{n\sigma}\subseteq S^{\prime n\sigma}$. Apply this to $S^{\prime}=R^{\nu}$, noting that $(R^{\nu})^{n\sigma}=R^{\nu}$ as $R^{\nu}$ is a Krull domain [Nagata-local, Theorem 33.10], we see $(i)$ holds.
Now apply the inclusion $S^{n\sigma}\subseteq S^{\prime n\sigma}$ and Lemma 4.3 to a sub-$S$-algebra $S^{\prime}$ of $S^{n\sigma}=S^{\sigma}$. We get $S^{n\sigma}=S^{\prime n\sigma}$, giving $(ii)$$(iii)$. ∎
@@ 5. Extension rings via first local cohomology
This is an auxiliary section that provides a conceptual variant of Ratliff’s construction [Mat-CRT, §31, Lemma 4].
@@
Let $M_{1}$ be the socle of $H^{1}_{\mathfrak{m}}(R)$, which is finite as $H^{1}_{\mathfrak{m}}(R)$ is Artinian [Residues-Duality, Chapter V, Corollary 6.5]. The ring $R_{1}:=R[\mathfrak{m};M_{1}]$ is then finite over $R$ by Lemma 5.5. As $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch, $R_{1}$ is a local ring, and we denote its maximal ideal by $\mathfrak{m}_{1}$. We know $\dim R_{1}=\dim R\geq 2$, and we know $0ptR_{1}\geq 1$ as $R_{1}$ is a subring of $\mathcal{O}(D(\mathfrak{m}))$. Therefore the same applies to the ring $R_{1}$, and from Discussion 5.3 we see $R[\mathfrak{m};M_{2}]$ is finite over $R$ for $M_{2}=(M_{1}^{a}:\mathfrak{m})$. Inductively we see $R[\mathfrak{m};H^{1}_{\mathfrak{m}}(R)]$ is integral over $R$, and so is its subring $R[\mathfrak{m};M]$. ∎
@@ 6. Eliminating the obstructions
The following result is [Mat-CRT, Theorem 31.1], and also follows from [EGA4_2, Proposition 6.10.6].
@@
Let $R$ be a Noetherian semilocal domain. Then there exists a finite sub-$R$-algebra $R^{\prime}$ of $R^{\sigma}$ such that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$.
@@
By Lemmas 6.1 and 6.3 it suffices to find a finite sub-$R$-algebra $R^{\prime}$ of $R^{\sigma}$ for each $P\in\operatorname{OS}_{2}^{\wedge}({R})$ such that no FONSIs of $R^{\prime}$ are above $P$. Fix a $P\in\operatorname{OS}_{2}^{\wedge}({R})$ and let $\mathfrak{p}=P\cap R\in\operatorname{Spec}(R)$, so we have $\operatorname{ht}(\mathfrak{p})>1$.
By Theorem 3.1, there exists a finite $D(\mathfrak{p})$-modification $R_{1}$ of $R$ such that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{1})$ are p-unibranch. We know $R_{1}\subseteq R^{\sigma}$ as $\operatorname{ht}(\mathfrak{p})>1$. Next, we take a finite $D(\mathfrak{p}R_{1})$-modification $R_{2}$ of $R_{1}$, so that the number of maximal ideals of $(R_{2}^{\wedge})_{P}$ (i.e. the number of preimages of $P$ in $\operatorname{Spec}(R_{2}^{\wedge})$) is maximal among all possible $R_{2}$; to see this is achievable, note that $(R_{2}^{\wedge})_{\operatorname{red}}$ is finite birational over $(R_{1}^{\wedge})_{\operatorname{red}}$, so $((R_{2}^{\wedge})_{\operatorname{red}})_{P}$ is contained in the normalization of $((R_{1}^{\wedge})_{\operatorname{red}})_{P}$, which is finite as a complete Noetherian local ring is Nagata [stacks, Tag 0335], so the number of maximal ideals of $(R_{2}^{\wedge})_{P}$ is bounded. Note that we still have $R_{2}\subseteq R^{\sigma}$ and that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{2})$ are p-unibranch. We will show no FONSIs of $R_{2}$ are above $P$.
@@
That $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$ implies $R^{n\sigma}$ being integral over $R$ is Theorem 4.8$(i)$. Assume $\operatorname{OS}_{2}^{\wedge}({R})\neq\emptyset$. We show $R^{n\sigma}$ is not integral over $R$. By Lemma 4.2 we may assume $R$ is local. Let $R^{\prime}$ be as in Theorem 6.4. Let $P\in\operatorname{OS}_{2}^{\wedge}({R})$ and let $P_{0}\in\operatorname{Min}(R^{\wedge})$ be contained in $P$ so that $\operatorname{ht}(P/P_{0})=1$. Write $\mathfrak{p}=P\cap R$, so $\operatorname{ht}(\mathfrak{p})>1$.
@@
To see the claim, let $\mathfrak{q}\in\operatorname{Spec}_{1}(R)$ be contained in $\mathfrak{p}$. As $\operatorname{ht}(\mathfrak{p})>1$, we can find $x\in\mathfrak{p}\setminus\mathfrak{q}$. Let $\beta\in W^{\prime-1}R^{\prime}$. Since $\operatorname{ht}(\mathfrak{p}^{\prime})=1$, we know the fraction field of $R^{\prime}_{\mathfrak{p}^{\prime}}$, which is also the fraction field of $R$ and $R^{\prime}$, is equal to $R^{\prime}_{\mathfrak{p}^{\prime}}[\frac{1}{x}]$. Therefore $x^{N}\beta\in R^{\prime}_{\mathfrak{p}^{\prime}}$ for some $N$, so $x^{N}\beta\in R^{\prime}_{\mathfrak{p}}$ as $\beta\in W^{\prime-1}R^{\prime}$. As $R^{\prime}\subseteq R^{\sigma}$, we have $R_{\mathfrak{q}}=R^{\prime}_{\mathfrak{q}}$. Thus $x^{N}\beta\in R_{\mathfrak{q}}$, so $\beta\in R_{\mathfrak{q}}$ as $x\not\in\mathfrak{q}$. ∎
@@
Let $R$ be a Noetherian integral domain. Then the canonical $(S_{2})$-ification $R^{\sigma}$ is $(S_{2})$.
@@
By Lemmas 4.2 and 3.6, we may assume $R$ is local. By Theorem 6.4 we can find a finite $R^{\prime}\subseteq R^{\sigma}$ so that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$. Let $\mathcal{R}$ be the set of all finite sub-$R^{\prime}$-algebras of $R^{\sigma}$. For every $R^{\prime\prime}\in\mathcal{R}$ we have $R^{\prime\prime\sigma}\subseteq R^{\sigma}$ by Lemma 4.3, so $R^{\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime\sigma}$, and this union is filtered by Lemmas 6.1 and 4.10. By Lemma 6.1 and Theorem 4.8, each $R^{\prime\prime\sigma}$ is $(S_{2})$. We conclude by Lemma 3.7. ∎
@@
Let $R$ be a Nagata integral domain. Then there exists a (finite) sub-$R$-algebra $R^{\prime}$ of $R^{\nu}$ that is $(S_{2})$ and satisfies $R_{\mathfrak{p}}=R^{\prime}_{\mathfrak{p}}$ for all $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$.
@@
Let $R$ be a Noetherian integral domain, $S$ a sub-$R$-algebra of $R^{\nu}$. Then $S^{\sigma}$ is $(S_{2})$.
@@
By Lemmas 4.2 and 3.6, we may assume $R$ is local. Let $\Sigma=\{\mathfrak{q}\in\operatorname{Spec}_{1}(S)\mid\operatorname{ht}(\mathfrak{q}\cap R)>1\}$. Then by Lemmas 4.10 and 6.3 and Theorem 3.1, $\Sigma$ is finite. We may therefore find a finite $R$-subalgebra $R^{\prime}$ of $S$ such that $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ for all $\mathfrak{q}\in\Sigma$, Lemma 3.2. For all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)\setminus\Sigma$, $\operatorname{ht}(\mathfrak{q}\cap R)=1$, so $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ as $R\to R^{\prime}$ is integral. Therefore $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ for all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)$, consequently $\operatorname{ht}(\mathfrak{q}\cap R^{\prime\prime})=1$ for all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)$ and all $R^{\prime\prime}\in\mathcal{R}:=$ the set of all finite sub-$R^{\prime}$-algebras of $S$. As in the proof of Corollary 4.11, we see $S^{n\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime n\sigma}$, so $S^{\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime\sigma}$ as $S^{\nu}=R^{\prime\prime\nu}=R^{\nu}$ for all $R^{\prime\prime}\in\mathcal{R}$. Again, as in the proof of Corollary 4.11, this union is filtered, and we conclude by Lemma 3.7 and Theorem 6.6. ∎
@@ 7. Semi-Nagata rings
Let $x\in R^{\circ}$ be a noninvertible element. Then $R^{\prime}/xR^{\prime}$ is $(S_{1})$. As the fibers of $R/xR\to(R/xR)^{\wedge}$ are $(S_{1})$ by the induction hypothesis, so are the fibers of $R^{\prime}/xR^{\prime}\to(R^{\prime}/xR^{\prime})^{\wedge}$. Therefore $(R^{\prime}/xR^{\prime})^{\wedge}$ is $(S_{1})$ [stacks, Tag 0339], so $R^{\prime\wedge}$ is $(S_{1})$ [EGA4_2, Proposition 3.4.4]. As $R\subseteq R^{\prime}$ is a finite inclusion of Noetherian semilocal domains we see $R^{\wedge}$ is $(S_{1})$ (Lemma 2.1). ∎
@@
Let $R$ be a Noetherian semilocal domain such that $R^{\wedge}$ is $(S_{1})$. Then there exists a finite sub-$R$-algebra of $R^{\sigma}$ that is $(S_{2})$.
@@
By Theorem 6.4 we can find a finite sub-$R$-algebra $R^{\prime}$ of $R^{\sigma}$ so that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$. $R^{\prime\wedge}$ is $(S_{1})$ by Lemma 2.1. Therefore $R^{\prime\sigma}$ is a finite $R^{\prime}$-algebra inside $R^{\sigma}$ (Lemma 4.3) that is $(S_{2})$ (Theorem 4.8). ∎
@@ 2. $(ii)$
    For every finite ring map $R\to B$ where $B$ is an integral domain, there exists a finite sub-$B$-algebra $C$ of $B^{\sigma}$ that is $(S_{2})$.
@@ 3. $(iii)$
It is possible that $R^{\sigma}$ is not finite over $R$, even when $R$ is semi-Nagata. As in Example 4.4, there is a Noetherian local domain $(R,\mathfrak{m})$ of dimension $2$ such that $R^{\wedge}\cong k[[x,y,z]]/(x^{2},y^{2})\cap(z)$, where $k$ is a field. Then $k[[x,y,z]]/(x^{2},y^{2})\times k[[x,y]]$ is a finite $D(\mathfrak{m}R^{\wedge})$-modification of $R^{\wedge}$, thus isomorphic to $R^{\prime\wedge}$ where $R^{\prime}$ is a finite $D(\mathfrak{m})$-modification of $R$ [stacks, Tags 0ALK and 05EU]. By [EGA4_2, Proposition 6.3.8] the Cohen–Macaulay ring $R^{\prime}$ is an $(S_{1})$-ring, so $R^{\prime}$ is semi-Nagata, thus so is $R$. On the other hand, for the maximal ideal $\mathfrak{m}^{\prime}$ of $R^{\prime}$ of height $1$, the normalization $T$ of $R^{\prime}_{\mathfrak{m}^{\prime}}$ is contained in a localization of $R^{\sigma}$, as $\operatorname{ht}(\mathfrak{m}^{\prime}\cap R)=2$. Since $T$ is not finite over $R^{\prime}_{\mathfrak{m}^{\prime}}$ (otherwise the nonreduced ring $R^{\prime\wedge}_{\mathfrak{m}^{\prime}}\cong k[[x,y,z]]/(x^{2},y^{2})$ will be a subring of a finite product of DVRs), we see $R^{\sigma}$ is not finite over $R$.
@@
It cannot be expected that $C$ as in Theorem 7.8$(ii)$ is $(S_{2})$ as a $B$-module. In fact, for the semi-Nagata ring $R$ in Example 4.4 (or 7.9) there exists no inclusion of finite $R$-modules $R\to M$ so that $M$ is $(S_{2})$. To see this, as $\dim R=2$ we have $0ptM\geq 2$ (so $0pt(M^{\wedge})\geq 2$). At this point, we can apply [stacks, Tag 00NM] to see $R$ universally catenary, so $R^{\wedge}$ is equidimensional [stacks, Tag 0AW6], which is a contradiction. For a more explicit examination of what failed, by [stacks, Tags 0AVZ and 0DWR] we see $M^{\wedge}=\Gamma(U^{\wedge},\mathcal{F}^{\wedge})$, where $U^{\wedge}$ is the punctured spectrum of $R^{\wedge}$ and $\mathcal{F}^{\wedge}$ is the sheaf on $\operatorname{Spec}(R^{\wedge})$ associated with $M^{\wedge}$. Then for the minimal prime $P_{0}$ of $R^{\wedge}$ with $\dim(R^{\wedge}/P_{0})=1$, we see $(M^{\wedge})_{P_{0}}$ is a direct factor of $M^{\wedge}$. Therefore $(M^{\wedge})_{P_{0}}$, and its submodule $(R^{\wedge})_{P_{0}}$, is finite over $R^{\wedge}$; in other words, $k((z))$ is finite over $k[[z]]$, contradiction.

### v2

@@ 3. $(iii)$
    If a subalgebra $S^{\prime}$ of $S^{\sigma}$ is $(S_{2})$, then $S^{\prime}=S^{\sigma}$.
@@
Let $S^{\prime}$ be a $S$-subalgebra of $R^{\nu}$. Let $\mathfrak{q}^{\prime}\in\operatorname{Spec}_{1}(S^{\prime})$. Then $\operatorname{ht}(\mathfrak{q}^{\prime}\cap R)=1$ by Lemma 4.10, so $\operatorname{ht}(\mathfrak{q}^{\prime}\cap S)=1$. Therefore $S^{n\sigma}\subseteq S^{\prime n\sigma}$. Apply this to $S^{\prime}=R^{\nu}$, noting that $(R^{\nu})^{n\sigma}=R^{\nu}$ as $R^{\nu}$ is a Krull domain [Nag62, Theorem 33.10], we see $(i)$ holds.
Now apply the inclusion $S^{n\sigma}\subseteq S^{\prime n\sigma}$ and Lemma 4.3 to a $S$-subalgebra $S^{\prime}$ of $S^{n\sigma}=S^{\sigma}$. We get $S^{n\sigma}=S^{\prime n\sigma}$, giving (ii)(ii)$(iii)$. ∎
@@ 5. Extension rings via first local cohomology
This is an auxiliary section that provides a conceptual variant of Ratliff’s construction [Mat89, §31, Lemma 4].
@@
Let $M_{1}$ be the socle of $H^{1}_{\mathfrak{m}}(R)$, which is finite as $H^{1}_{\mathfrak{m}}(R)$ is Artinian [Har66, Chapter V, Corollary 6.5]. The ring $R_{1}:=R[\mathfrak{m};M_{1}]$ is then finite over $R$ by Lemma 5.5. As $\mathfrak{m}\in\operatorname{Spec}(R)$ is p-unibranch, $R_{1}$ is a local ring, and we denote its maximal ideal by $\mathfrak{m}_{1}$. We know $\dim R_{1}=\dim R\geq 2$, and we know $0ptR_{1}\geq 1$ as $R_{1}$ is a subring of $\mathcal{O}(D(\mathfrak{m}))$. Therefore the same applies to the ring $R_{1}$, and from Discussion 5.3 we see $R[\mathfrak{m};M_{2}]$ is finite over $R$ for $M_{2}=(M_{1}^{a}:\mathfrak{m})$. Inductively we see $R[\mathfrak{m};H^{1}_{\mathfrak{m}}(R)]$ is integral over $R$, and so is its subring $R[\mathfrak{m};M]$. ∎
@@ 6. Eliminating the obstructions
The following result is [Mat89, Theorem 31.1], and also follows from [EGA ̵IV2, Proposition 6.10.6].
@@
Let $R$ be a Noetherian semilocal domain. Then there exists a finite subalgebra $R^{\prime}$ of $R^{\sigma}$ such that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$.
@@
By Lemmas 6.1 and 6.3 it suffices to find a finite subalgebra $R^{\prime}$ of $R^{\sigma}$ for each $P\in\operatorname{OS}_{2}^{\wedge}({R})$ such that no FONSIs of $R^{\prime}$ are above $P$. Fix a $P\in\operatorname{OS}_{2}^{\wedge}({R})$ and let $\mathfrak{p}=P\cap R\in\operatorname{Spec}(R)$, so we have $\operatorname{ht}(\mathfrak{p})>1$.
By Theorem 3.1, there exists a finite $D(\mathfrak{p})$-modification $R_{1}$ of $R$ such that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{1})$ are p-unibranch. We know $R_{1}\subseteq R^{\sigma}$ as $\operatorname{ht}(\mathfrak{p})>1$. Next, we take a finite $D(\mathfrak{p}R_{1})$-modification $R_{2}$ of $R_{1}$, so that the number of maximal ideals of $(R_{2}^{\wedge})_{\mathfrak{P}}$ (i.e. the number of preimages of $P$ in $\operatorname{Spec}(R_{2}^{\wedge})$) is maximal among all possible $R_{2}$; to see this is achievable, note that $(R_{2}^{\wedge})_{\operatorname{red}}$ is finite birational over $(R_{1}^{\wedge})_{\operatorname{red}}$, so $((R_{2}^{\wedge})_{\operatorname{red}})_{\mathfrak{P}}$ is contained in the normalization of $((R_{1}^{\wedge})_{\operatorname{red}})_{\mathfrak{P}}$, which is finite as a complete Noetherian local ring is Nagata [Stacks, Tag 0335], so the number of maximal ideals of $(R_{2}^{\wedge})_{\mathfrak{P}}$ is bounded. Note that we still have $R_{2}\subseteq R^{\sigma}$ and that all preimages of $\mathfrak{p}$ in $\operatorname{Spec}(R_{2})$ are p-unibranch. We will show no FONSIs of $R_{2}$ are above $P$.
@@
That $\operatorname{OS}_{2}^{\wedge}({R})=\emptyset$ implies $R^{n\sigma}$ being integral over $R$ is Theorem 4.8(i)(i). Assume $\operatorname{OS}_{2}^{\wedge}({R})\neq\emptyset$. We show $R^{n\sigma}$ is not integral over $R$. By Lemma 4.2 we may assume $R$ is local. Let $R^{\prime}$ be as in Theorem 6.4. Let $P\in\operatorname{OS}_{2}^{\wedge}({R})$ and let $P_{0}\in\operatorname{Min}(R^{\wedge})$ be contained in $P$ so that $\operatorname{ht}(P/P_{0})=1$. Write $\mathfrak{p}=P\cap R$, so $\operatorname{ht}(\mathfrak{p})>1$.
@@
To see the claim, let $\mathfrak{q}\in\operatorname{Spec}_{1}(R)$ be contained in $\mathfrak{p}$. We need to show $W^{\prime-1}R^{\prime}\subseteq R_{\mathfrak{q}}$. As $\operatorname{ht}(\mathfrak{p})>1$, we can find $x\in\mathfrak{p}\setminus\mathfrak{q}$. Let $\beta\in W^{\prime-1}R^{\prime}$. Since $\operatorname{ht}(\mathfrak{p}^{\prime})=1$, we know the fraction field of $R^{\prime}_{\mathfrak{p}^{\prime}}$, which is also the fraction field of $R$ and $R^{\prime}$, is equal to $R^{\prime}_{\mathfrak{p}^{\prime}}[\frac{1}{x}]$. Therefore $x^{N}\beta\in R^{\prime}_{\mathfrak{p}^{\prime}}$ for some $N$, so $x^{N}\beta\in R^{\prime}_{\mathfrak{p}}$ as $\beta\in W^{\prime-1}R^{\prime}$. As $R^{\prime}\subseteq R^{\sigma}$, we have $R_{\mathfrak{q}}=R^{\prime}_{\mathfrak{q}}$. Thus $x^{N}\beta\in R_{\mathfrak{q}}$, so $\beta\in R_{\mathfrak{q}}$ as $x\not\in\mathfrak{q}$. ∎
@@
Let $R$ be a Noetherian integral domain. Then $R^{\sigma}$ is $(S_{2})$.
@@
By Lemmas 4.2 and 3.6, we may assume $R$ is local. By Theorem 6.4 we can find a finite $R^{\prime}\subseteq R^{\sigma}$ so that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$. Let $\mathcal{R}$ be the set of all finite $R^{\prime}$-subalgebras of $R^{\sigma}$. For every $R^{\prime\prime}\in\mathcal{R}$ we have $R^{\prime\prime\sigma}\subseteq R^{\sigma}$ by Lemma 4.3, so $R^{\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime\sigma}$, and this union is filtered by Lemmas 6.1 and 4.10. By Lemma 6.1 and Theorem 4.8, each $R^{\prime\prime\sigma}$ is $(S_{2})$. We conclude by Lemma 3.7. ∎
@@
Let $R$ be a Nagata integral domain. Then there exists a (finite) subalgebra $R^{\prime}$ of $R^{\nu}$ that is $(S_{2})$ and satisfies $R_{\mathfrak{p}}=R^{\prime}_{\mathfrak{p}}$ for all $\mathfrak{p}\in\operatorname{Spec}_{1}(R)$.
@@
Let $R$ be a Noetherian integral domain, $S$ a subalgebra of $R^{\nu}$. Then $S^{\sigma}$ is $(S_{2})$.
@@
By Lemmas 4.2 and 3.6, we may assume $R$ is local. Let $\Sigma=\{\mathfrak{q}\in\operatorname{Spec}_{1}(S)\mid\operatorname{ht}(\mathfrak{q}\cap R)>1\}$. Then by Lemmas 4.10 and 6.3 and Theorem 3.1, $\Sigma$ is finite. We may therefore find a finite subalgebra $R^{\prime}$ of $S$ such that $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ for all $\mathfrak{q}\in\Sigma$, Lemma 3.2. For all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)\setminus\Sigma$, $\operatorname{ht}(\mathfrak{q}\cap R)=1$, so $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ as $R\to R^{\prime}$ is integral. Therefore $\operatorname{ht}(\mathfrak{q}\cap R^{\prime})=1$ for all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)$, consequently $\operatorname{ht}(\mathfrak{q}\cap R^{\prime\prime})=1$ for all $\mathfrak{q}\in\operatorname{Spec}_{1}(S)$ and all $R^{\prime\prime}\in\mathcal{R}:=$ the set of all finite $R^{\prime}$-subalgebras of $S$. As in the proof of Corollary 4.11, we see $S^{n\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime n\sigma}$, so $S^{\sigma}=\bigcup_{R^{\prime\prime}\in\mathcal{R}}R^{\prime\prime\sigma}$ as $S^{\nu}=R^{\prime\prime\nu}=R^{\nu}$ for all $R^{\prime\prime}\in\mathcal{R}$. Again, as in the proof of Corollary 4.11, this union is filtered, and we conclude by Lemma 3.7 and Theorem 6.6. ∎
@@ 7. Semi-Nagata rings
Let $x\in R^{\circ}$ be a noninvertible element. Then $R^{\prime}/xR^{\prime}$ is $(S_{1})$. As the fibers of $R/xR\to(R/xR)^{\wedge}$ are $(S_{1})$ by the induction hypothesis, so are the fibers of $R^{\prime}/xR^{\prime}\to(R^{\prime}/xR^{\prime})^{\wedge}$. Therefore $(R^{\prime}/xR^{\prime})^{\wedge}$ is $(S_{1})$ [Stacks, Tag 0339], so $R^{\prime\wedge}$ is $(S_{1})$ [EGA ̵IV2, Proposition 3.4.4]. As $R\subseteq R^{\prime}$ is a finite inclusion of Noetherian semilocal domains we see $R^{\wedge}$ is $(S_{1})$ (Lemma 2.1). ∎
@@
Let $R$ be a Noetherian semilocal domain such that $R^{\wedge}$ is $(S_{1})$. Then there exists a finite subalgebra of $R^{\sigma}$ that is $(S_{2})$.
@@
By Theorem 6.4 we can find a finite subalgebra $R^{\prime}$ of $R^{\sigma}$ so that $\operatorname{OS}_{2}^{\wedge}({R^{\prime}})=\emptyset$. $R^{\prime\wedge}$ is $(S_{1})$ by Lemma 2.1. Therefore $R^{\prime\sigma}$ is a finite $R^{\prime}$-algebra inside $R^{\sigma}$ (Lemma 4.3) that is $(S_{2})$ (Theorem 4.8). ∎
@@ 2. $(ii)$
    For every finite ring map $R\to B$ where $B$ is an integral domain, there exists a finite subalgebra $C$ of $B^{\sigma}$ that is $(S_{2})$.
@@ 3. $(iii)$
It is possible that $R^{\sigma}$ is not finite over $R$, even when $R$ is semi-Nagata. As in Example 4.4, there is a Noetherian local domain $(R,\mathfrak{m})$ of dimension $2$ such that $R^{\wedge}\cong k[[x,y,z]]/(x^{2},y^{2})\cap(z)$, where $k$ is a field. Then $k[[x,y,z]]/(x^{2},y^{2})\times k[[x,y]]$ is a finite $D(\mathfrak{m}R^{\wedge})$-modification of $R^{\wedge}$, thus isomorphic to $R^{\prime\wedge}$ where $R^{\prime}$ is a finite $D(\mathfrak{m})$-modification of $R$ [Stacks, Tags 0ALK and 05EU]. By [EGA ̵IV2, Proposition 6.3.8] the Cohen–Macaulay ring $R^{\prime}$ is an $(S_{1})$-ring, so $R^{\prime}$ is semi-Nagata, thus so is $R$. On the other hand, for the maximal ideal $\mathfrak{m}^{\prime}$ of $R^{\prime}$ of height $1$, the normalization $T$ of $R^{\prime}_{\mathfrak{m}^{\prime}}$ is contained in a localization of $R^{\sigma}$, as $\operatorname{ht}(\mathfrak{m}^{\prime}\cap R)=2$. Since $T$ is not finite over $R^{\prime}_{\mathfrak{m}^{\prime}}$ (otherwise the nonreduced ring $R^{\prime\wedge}_{\mathfrak{m}^{\prime}}\cong k[[x,y,z]]/(x^{2},y^{2})$ will be a subring of a finite product of DVRs), we see $R^{\sigma}$ is not finite over $R$.
@@
In general, it cannot be expected that $C$ as in Theorem 7.8(ii)(ii) is $(S_{2})$ as a $B$-module. In fact, for the semi-Nagata ring $R$ in Example 4.4 (or 7.9) there exists no inclusion of finite $R$-modules $R\to M$ so that $M$ is $(S_{2})$. To see this, as $\dim R=2$ we have $0ptM\geq 2$ (so $0pt(M^{\wedge})\geq 2$). At this point, we can apply [Stacks, Tag 00NM] to see $R$ universally catenary, so $R^{\wedge}$ is equidimensional [Stacks, Tag 0AW6], which is a contradiction. For a more explicit examination of what failed, by [Stacks, Tags 0AVZ and 0DWR] we see $M^{\wedge}=\Gamma(U^{\wedge},\mathcal{F}^{\wedge})$, where $U^{\wedge}$ is the punctured spectrum of $R^{\wedge}$ and $\mathcal{F}^{\wedge}$ is the sheaf on $\operatorname{Spec}(R^{\wedge})$ associated with $M^{\wedge}$. Then for the minimal prime $P_{0}$ of $R^{\wedge}$ with $\dim(R^{\wedge}/P_{0})=1$, we see $(M^{\wedge})_{P_{0}}$ is a direct factor of $M^{\wedge}$. Therefore $(M^{\wedge})_{P_{0}}$, and its submodule $(R^{\wedge})_{P_{0}}$, is finite over $R^{\wedge}$; in other words, $k((z))$ is finite over $k[[z]]$, contradiction.
@@
Remark 7.11.
On the other hand, if $R$ is a Noetherian universally catenary domain and $R^{\prime}$ is an integral domain containing and finite over $R$, then $R^{\prime}$ is $(S_{2})$ as a ring if and only if $R^{\prime}$ as an $R$-module. To see this, let us assume $(R,\mathfrak{m})$ local, so $R^{\prime}$ is semilocal with maximal ideals $\mathfrak{m}^{\prime}_{1},\ldots,\mathfrak{m}^{\prime}_{n}$. Since $R$ is universally catenary we know $\operatorname{ht}(\mathfrak{m}^{\prime}_{i})=\dim R$ for every $i$, see [Stacks, Tag 02IJ]. Moreover, as $\mathfrak{m}^{\prime}_{i}$ are the only preimages of $\mathfrak{m}$ in $\operatorname{Spec}(R^{\prime})$, we have $\min_{i}0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}=0pt_{R}R^{\prime}$. We conclude that $0pt_{R}R^{\prime}\geq\max\{2,\dim R\}$ if and only if $0ptR^{\prime}_{\mathfrak{m}^{\prime}_{i}}\geq\max\{2,\operatorname{ht}(\mathfrak{m}^{\prime}_{i})\}$ for all $i$. This gives the original assertion via localization.